# Prompt Optimization with EASE? Efficient Ordering-aware Automated Selection of Exemplars

### v1

@@ C.10 Test accuracies for all tables
antonyms 75.7$\pm$0.3 86.0$\pm$0.0 85.3$\pm$0.5 77.7$\pm$0.3 80.7$\pm$0.3 80.0$\pm$2.5 82.3$\pm$0.3 82.0$\pm$0.0 84.3$\pm$0.3 81.0$\pm$0.0
auto_categorization 28.7$\pm$0.3 25.7$\pm$0.5 24.0$\pm$0.5 39.3$\pm$0.7 27.0$\pm$0.8 32.7$\pm$1.0 35.0$\pm$1.2 17.3$\pm$0.3 30.7$\pm$0.5 31.3$\pm$0.5
@@
object_counting 45.3$\pm$0.7 48.3$\pm$0.7 40.7$\pm$0.3 27.0$\pm$2.0 31.3$\pm$3.1 48.0$\pm$2.0 58.0$\pm$0.5 42.3$\pm$0.3 52.3$\pm$2.7 55.3$\pm$5.2
orthography_starts_with 30.3$\pm$0.7 42.0$\pm$0.5 65.7$\pm$0.3 65.7$\pm$0.7 71.0$\pm$0.5 40.3$\pm$7.8 69.3$\pm$0.7 47.0$\pm$0.8 63.3$\pm$2.0 67.3$\pm$1.1
rhymes 58.0$\pm$0.8 42.0$\pm$0.5 11.3$\pm$0.7 99.0$\pm$0.0 64.3$\pm$1.5 47.3$\pm$6.9 59.3$\pm$12.6 98.3$\pm$0.3 96.0$\pm$0.0 89.7$\pm$1.9
second_word_letter 17.3$\pm$0.7 31.0$\pm$0.0 32.3$\pm$0.5 42.0$\pm$2.5 42.3$\pm$0.7 27.0$\pm$8.2 45.7$\pm$1.0 40.3$\pm$0.3 38.7$\pm$0.3 42.0$\pm$0.5
sentence_similarity 19.0$\pm$0.5 13.7$\pm$0.3 38.3$\pm$1.2 41.0$\pm$0.5 33.7$\pm$1.9 19.0$\pm$1.7 20.0$\pm$4.1 18.0$\pm$0.0 31.3$\pm$2.2 30.0$\pm$2.2
sentiment 91.3$\pm$0.3 89.0$\pm$0.0 87.0$\pm$0.0 90.3$\pm$0.3 90.7$\pm$0.5 91.7$\pm$0.7 89.0$\pm$0.0 89.7$\pm$0.3 89.3$\pm$0.3 88.0$\pm$0.8
@@
synonyms 14.0$\pm$0.5 14.3$\pm$0.5 11.0$\pm$0.0 14.0$\pm$0.0 13.7$\pm$0.3 13.0$\pm$0.5 14.7$\pm$0.7 13.3$\pm$0.7 18.3$\pm$0.3 14.0$\pm$1.2
taxonomy_animal 47.7$\pm$0.3 44.7$\pm$1.0 77.3$\pm$1.5 73.3$\pm$0.3 67.3$\pm$4.8 46.3$\pm$7.1 62.3$\pm$7.3 32.7$\pm$0.5 73.7$\pm$3.4 74.7$\pm$5.5
translation_en-de 83.7$\pm$0.3 84.3$\pm$0.3 83.3$\pm$0.3 83.0$\pm$0.5 83.3$\pm$1.1 80.3$\pm$1.9 85.0$\pm$0.9 83.3$\pm$0.3 82.3$\pm$0.3 83.7$\pm$0.5
translation_en-es 83.7$\pm$0.3 89.7$\pm$0.3 88.3$\pm$0.3 87.3$\pm$1.0 87.3$\pm$1.0 88.0$\pm$0.8 86.7$\pm$1.2 84.3$\pm$0.3 84.3$\pm$0.3 85.0$\pm$0.0
translation_en-fr 83.0$\pm$0.0 87.7$\pm$0.3 87.3$\pm$0.3 87.0$\pm$0.0 87.0$\pm$0.0 83.3$\pm$1.7 87.7$\pm$0.5 86.7$\pm$0.3 88.3$\pm$0.3 87.3$\pm$0.7
word_sorting 8.7$\pm$0.5 65.0$\pm$0.5 66.7$\pm$1.0 66.0$\pm$0.9 43.7$\pm$1.9 71.0$\pm$1.4 69.3$\pm$1.4 61.3$\pm$1.0 63.3$\pm$0.5 68.0$\pm$2.0
word_unscrambling 61.3$\pm$0.3 58.7$\pm$0.5 63.3$\pm$0.3 57.0$\pm$0.9 61.3$\pm$0.5 60.7$\pm$1.7 58.3$\pm$1.2 53.7$\pm$0.3 60.3$\pm$0.3 61.7$\pm$1.0
# best-performing tasks 4 6 7 8 5 7 9 7 8 7
@@
Rule-based tasks LR 0% 39.3$\pm$0.5 35.0$\pm$0.5 34.7$\pm$0.5 47.3$\pm$0.3 36.3$\pm$1.2 34.0$\pm$3.7 53.3$\pm$1.1 38.3$\pm$0.7 51.7$\pm$2.7 49.3$\pm$2.0
10% 0.0$\pm$0.0 37.0$\pm$0.5 34.0$\pm$0.5 49.0$\pm$0.9 39.3$\pm$1.8 39.0$\pm$0.8 52.3$\pm$2.1 37.7$\pm$0.5 48.0$\pm$0.9 46.0$\pm$0.8
30% 0.0$\pm$0.0 44.3$\pm$0.5 36.0$\pm$0.9 38.7$\pm$3.5 34.3$\pm$1.5 36.7$\pm$2.4 48.0$\pm$2.5 35.7$\pm$0.3 49.3$\pm$0.5 49.3$\pm$1.1
50% 0.0$\pm$0.0 53.3$\pm$1.0 31.0$\pm$0.5 47.0$\pm$0.5 34.3$\pm$1.1 12.7$\pm$5.2 42.7$\pm$1.7 34.0$\pm$0.5 49.3$\pm$0.3 54.3$\pm$2.1
70% 0.0$\pm$0.0 39.0$\pm$1.2 30.3$\pm$0.7 46.7$\pm$1.1 36.3$\pm$0.5 4.3$\pm$3.5 51.0$\pm$1.4 33.3$\pm$0.5 31.3$\pm$0.7 49.0$\pm$3.3
90% 0.0$\pm$0.0 35.0$\pm$0.5 39.7$\pm$0.3 35.3$\pm$1.5 0.0$\pm$0.0 0.0$\pm$0.0 41.0$\pm$1.4 16.7$\pm$0.7 28.0$\pm$0.5 41.3$\pm$1.9
LP-variant 0% 39.3$\pm$0.5 35.0$\pm$0.5 34.7$\pm$0.5 47.3$\pm$0.3 36.3$\pm$1.2 34.0$\pm$3.7 53.3$\pm$1.1 38.3$\pm$0.7 51.7$\pm$2.7 49.3$\pm$2.0
10% 0.0$\pm$0.0 37.0$\pm$0.5 34.0$\pm$0.5 49.0$\pm$0.9 39.3$\pm$1.8 39.0$\pm$0.8 52.3$\pm$2.1 37.7$\pm$0.5 48.0$\pm$0.9 46.0$\pm$0.8
30% 0.0$\pm$0.0 44.3$\pm$0.5 36.0$\pm$0.9 38.7$\pm$3.5 34.3$\pm$1.5 36.7$\pm$2.4 48.0$\pm$2.5 35.7$\pm$0.3 49.3$\pm$0.5 49.3$\pm$1.1
50% 0.0$\pm$0.0 53.3$\pm$1.0 31.0$\pm$0.5 47.0$\pm$0.5 34.3$\pm$1.1 12.7$\pm$5.2 42.7$\pm$1.7 34.0$\pm$0.5 49.3$\pm$0.3 54.3$\pm$2.1
70% 0.0$\pm$0.0 39.0$\pm$1.2 30.3$\pm$0.7 46.7$\pm$1.1 36.3$\pm$0.5 4.3$\pm$3.5 51.0$\pm$1.4 33.3$\pm$0.5 31.3$\pm$0.7 49.0$\pm$3.3
90% 0.0$\pm$0.0 35.0$\pm$0.5 39.7$\pm$0.3 35.3$\pm$1.5 0.0$\pm$0.0 0.0$\pm$0.0 41.0$\pm$1.4 16.7$\pm$0.7 28.0$\pm$0.5 41.3$\pm$1.9
Re-mapped label tasks AG News Remap 0% 7.0$\pm$0.5 7.0$\pm$0.0 12.3$\pm$0.5 30.3$\pm$1.2 28.3$\pm$1.1 4.3$\pm$1.1 9.0$\pm$2.0 11.3$\pm$0.3 19.3$\pm$0.5 29.7$\pm$1.9
10% 2.0$\pm$0.0 7.0$\pm$0.0 13.0$\pm$0.9 17.0$\pm$3.3 19.0$\pm$3.7 6.0$\pm$0.8 13.0$\pm$2.6 12.0$\pm$0.5 30.0$\pm$0.5 36.3$\pm$1.0
30% 3.7$\pm$0.3 1.0$\pm$0.0 4.7$\pm$0.3 28.0$\pm$1.6 22.3$\pm$0.5 7.7$\pm$2.2 5.0$\pm$0.5 12.3$\pm$0.3 28.0$\pm$4.1 32.3$\pm$3.7
50% 4.7$\pm$0.3 3.0$\pm$0.0 3.7$\pm$0.3 27.3$\pm$1.9 21.3$\pm$2.2 8.7$\pm$3.1 13.0$\pm$2.9 6.7$\pm$0.3 21.7$\pm$1.9 43.7$\pm$2.8
70% 2.0$\pm$0.0 23.3$\pm$0.3 7.0$\pm$0.0 20.0$\pm$0.8 16.0$\pm$0.5 15.7$\pm$7.1 5.0$\pm$0.9 4.7$\pm$0.3 36.0$\pm$0.5 31.0$\pm$2.5
90% 8.0$\pm$0.5 21.0$\pm$0.9 2.0$\pm$0.0 22.7$\pm$5.7 2.0$\pm$0.0 11.7$\pm$1.9 23.0$\pm$1.2 6.3$\pm$0.3 27.3$\pm$1.5 40.7$\pm$3.1
SST5 Reverse 0% 13.3$\pm$0.5 11.7$\pm$0.5 9.7$\pm$0.7 22.0$\pm$2.9 18.3$\pm$0.5 10.3$\pm$0.3 21.7$\pm$4.5 11.3$\pm$0.3 23.7$\pm$0.5 32.3$\pm$2.8
10% 13.3$\pm$0.7 11.3$\pm$0.3 9.3$\pm$0.3 20.0$\pm$2.2 18.0$\pm$0.8 11.7$\pm$1.5 27.0$\pm$0.9 12.0$\pm$0.5 23.0$\pm$0.5 30.7$\pm$1.9
30% 15.7$\pm$0.3 13.3$\pm$0.3 11.7$\pm$0.5 19.3$\pm$1.9 23.7$\pm$0.5 9.7$\pm$0.3 21.3$\pm$2.2 11.0$\pm$0.5 13.0$\pm$1.2 26.3$\pm$2.2
50% 13.0$\pm$0.5 12.3$\pm$0.3 12.0$\pm$0.0 28.3$\pm$2.4 14.0$\pm$0.5 12.7$\pm$0.7 14.3$\pm$0.7 9.3$\pm$0.3 14.0$\pm$1.6 29.0$\pm$2.9
70% 12.0$\pm$0.0 12.0$\pm$0.5 14.3$\pm$0.3 24.7$\pm$0.5 12.0$\pm$0.9 11.7$\pm$0.5 18.7$\pm$2.3 15.7$\pm$0.3 33.0$\pm$0.8 33.0$\pm$1.4
90% 16.0$\pm$0.5 9.3$\pm$0.5 16.0$\pm$0.5 12.7$\pm$0.3 13.3$\pm$0.7 12.7$\pm$0.3 14.0$\pm$0.5 11.7$\pm$0.5 10.7$\pm$0.3 11.3$\pm$1.2
# best-performing tasks 1 0 1 1 0 0 6 0 4 13
@@
antonyms 81.0$\pm$0.0 80.0$\pm$1.2 -1.0 $\downarrow$
auto_categorization 31.3$\pm$0.5 35.3$\pm$4.5 4.0 $\uparrow$
@@
larger_animal 87.7$\pm$1.5 62.0$\pm$0.5 -25.7 $\downarrow$
negation 89.0$\pm$0.5 86.7$\pm$0.7 -2.3 $\downarrow$
object_counting 55.3$\pm$5.2 61.3$\pm$1.4 6.0 $\uparrow$
orthography_starts_with 67.3$\pm$1.1 72.3$\pm$1.4 5.0 $\uparrow$
rhymes 89.7$\pm$1.9 65.7$\pm$2.7 -24.0 $\downarrow$
second_word_letter 42.0$\pm$0.5 100.0$\pm$0.0 58.0 $\uparrow$
sentence_similarity 30.0$\pm$2.2 33.7$\pm$2.6 3.7 $\uparrow$
sentiment 88.0$\pm$0.8 92.3$\pm$0.5 4.3 $\uparrow$
@@
synonyms 14.0$\pm$1.2 14.3$\pm$0.7 0.3 $\uparrow$
taxonomy_animal 74.7$\pm$5.5 75.0$\pm$2.5 0.3 $\uparrow$
translation_en-de 83.7$\pm$0.5 83.0$\pm$0.0 -0.7 $\downarrow$
translation_en-es 85.0$\pm$0.0 88.7$\pm$0.7 3.7 $\uparrow$
translation_en-fr 87.3$\pm$0.7 87.3$\pm$0.5 0.0 $\circ$
word_sorting 68.0$\pm$2.0 73.7$\pm$0.7 5.7 $\uparrow$
word_unscrambling 61.7$\pm$1.0 60.7$\pm$0.3 -1.0 $\downarrow$
linear_4_10_noisy 48.3$\pm$1.5 27.0$\pm$10.5 -21.3 $\downarrow$
pig_latin_variant2_encode_10_noisy 46.0$\pm$0.8 57.3$\pm$1.0 11.3 $\uparrow$
ag_news_textlabel_redirect1_10_noisy 36.3$\pm$1.0 41.0$\pm$2.4 4.7 $\uparrow$
sst5_reverse_10_noisy 30.7$\pm$1.9 35.0$\pm$0.8 4.3 $\uparrow$

### v2

@@ C.15 Test accuracies for all tables
antonyms 75.7$\pm$0.3 86.0$\pm$0.0 85.3$\pm$0.7 77.7$\pm$0.3 80.7$\pm$0.3 80.0$\pm$2.5 82.3$\pm$0.3 82.0$\pm$0.0 84.3$\pm$0.3 84.7$\pm$0.3
auto_categorization 28.7$\pm$0.3 25.7$\pm$0.7 24.0$\pm$0.6 39.3$\pm$0.9 27.0$\pm$1.0 32.7$\pm$1.0 35.0$\pm$1.5 17.3$\pm$0.3 30.7$\pm$0.7 34.0$\pm$2.1
@@
object_counting 45.3$\pm$0.9 48.3$\pm$0.9 40.7$\pm$0.3 27.0$\pm$2.5 31.3$\pm$3.8 48.0$\pm$2.0 58.0$\pm$0.6 42.3$\pm$0.3 52.3$\pm$3.3 52.0$\pm$2.1
orthography_starts_with 30.3$\pm$0.9 42.0$\pm$0.6 65.7$\pm$0.3 65.7$\pm$0.9 71.0$\pm$0.6 40.3$\pm$7.8 69.3$\pm$0.9 47.0$\pm$1.0 63.3$\pm$2.4 69.3$\pm$0.7
rhymes 58.0$\pm$1.0 42.0$\pm$0.6 11.3$\pm$0.9 99.0$\pm$0.0 64.3$\pm$1.9 47.3$\pm$6.9 59.3$\pm$15.4 98.3$\pm$0.3 96.0$\pm$0.0 97.7$\pm$1.3
second_word_letter 17.3$\pm$0.9 31.0$\pm$0.0 32.3$\pm$0.7 42.0$\pm$3.1 42.3$\pm$0.9 27.0$\pm$8.2 45.7$\pm$1.2 40.3$\pm$0.3 38.7$\pm$0.3 48.3$\pm$1.4
sentence_similarity 19.0$\pm$0.6 13.7$\pm$0.3 38.3$\pm$1.4 41.0$\pm$0.6 33.7$\pm$2.3 19.0$\pm$1.7 20.0$\pm$5.0 18.0$\pm$0.0 31.3$\pm$2.7 32.7$\pm$1.2
sentiment 91.3$\pm$0.3 89.0$\pm$0.0 87.0$\pm$0.0 90.3$\pm$0.3 90.7$\pm$0.7 91.7$\pm$0.7 89.0$\pm$0.0 89.7$\pm$0.3 89.3$\pm$0.3 89.0$\pm$0.0
@@
synonyms 14.0$\pm$0.6 14.3$\pm$0.7 11.0$\pm$0.0 14.0$\pm$0.0 13.7$\pm$0.3 13.0$\pm$0.5 14.7$\pm$0.9 13.3$\pm$0.9 18.3$\pm$0.3 13.7$\pm$0.3
taxonomy_animal 47.7$\pm$0.3 44.7$\pm$1.2 77.3$\pm$1.9 73.3$\pm$0.3 67.3$\pm$5.9 46.3$\pm$7.1 62.3$\pm$8.9 32.7$\pm$0.7 73.7$\pm$4.2 78.7$\pm$1.9
translation_en-de 83.7$\pm$0.3 84.3$\pm$0.3 83.3$\pm$0.3 83.0$\pm$0.6 83.3$\pm$1.3 80.3$\pm$1.9 85.0$\pm$1.1 83.3$\pm$0.3 82.3$\pm$0.3 83.3$\pm$0.3
translation_en-es 83.7$\pm$0.3 89.7$\pm$0.3 88.3$\pm$0.3 87.3$\pm$1.2 87.3$\pm$1.2 88.0$\pm$0.8 86.7$\pm$1.4 84.3$\pm$0.3 84.3$\pm$0.3 84.7$\pm$0.9
translation_en-fr 83.0$\pm$0.0 87.7$\pm$0.3 87.3$\pm$0.3 87.0$\pm$0.0 87.0$\pm$0.0 83.3$\pm$1.7 87.7$\pm$0.7 86.7$\pm$0.3 88.3$\pm$0.3 85.0$\pm$1.0
word_sorting 8.7$\pm$0.7 65.0$\pm$0.6 66.7$\pm$1.2 66.0$\pm$1.1 43.7$\pm$2.3 71.0$\pm$1.4 69.3$\pm$1.7 61.3$\pm$1.2 63.3$\pm$0.7 69.7$\pm$0.3
word_unscrambling 61.3$\pm$0.3 58.7$\pm$0.7 63.3$\pm$0.3 57.0$\pm$1.1 61.3$\pm$0.7 60.7$\pm$1.7 58.3$\pm$1.4 53.7$\pm$0.3 60.3$\pm$0.3 62.0$\pm$2.1
# best-performing tasks 4 6 6 8 5 7 8 7 8 9
@@
Rule-based tasks LR 0% 36.7$\pm$0.7 38.0$\pm$1.0 34.0$\pm$1.0 48.0$\pm$1.1 40.7$\pm$2.3 32.0$\pm$6.6 43.3$\pm$5.8 38.0$\pm$0.6 47.7$\pm$0.3 58.0$\pm$2.1
10% 7.3$\pm$0.9 33.3$\pm$0.7 29.3$\pm$0.3 40.3$\pm$3.8 41.0$\pm$2.9 0.0$\pm$0.0 46.3$\pm$1.2 38.3$\pm$0.3 47.7$\pm$0.7 42.3$\pm$0.9
30% 12.0$\pm$0.6 23.3$\pm$0.3 30.0$\pm$1.0 45.7$\pm$4.8 29.7$\pm$2.9 9.0$\pm$5.8 35.0$\pm$4.7 38.3$\pm$1.2 43.0$\pm$0.6 47.7$\pm$5.8
50% 3.0$\pm$1.0 33.7$\pm$0.3 23.0$\pm$0.6 43.7$\pm$2.3 33.3$\pm$0.7 0.0$\pm$0.0 47.0$\pm$2.1 32.7$\pm$0.9 30.3$\pm$1.8 50.7$\pm$2.2
70% 0.0$\pm$0.0 38.7$\pm$0.3 28.7$\pm$0.7 44.0$\pm$3.5 34.0$\pm$1.1 1.7$\pm$1.4 31.7$\pm$3.8 30.3$\pm$0.7 26.3$\pm$5.3 47.0$\pm$3.2
90% 0.0$\pm$0.0 32.7$\pm$1.3 32.0$\pm$0.6 37.3$\pm$4.8 2.3$\pm$0.9 0.0$\pm$0.0 7.7$\pm$1.9 5.0$\pm$0.6 14.0$\pm$0.6 39.0$\pm$7.8
LP-variant 0% 39.3$\pm$0.7 35.0$\pm$0.6 34.7$\pm$0.7 47.3$\pm$0.3 36.3$\pm$1.4 34.0$\pm$3.7 53.3$\pm$1.3 38.3$\pm$0.9 51.7$\pm$3.3 53.0$\pm$2.3
10% 0.0$\pm$0.0 37.0$\pm$0.6 34.0$\pm$0.6 49.0$\pm$1.1 39.3$\pm$2.2 39.0$\pm$0.8 52.3$\pm$2.6 37.7$\pm$0.7 48.0$\pm$1.1 48.0$\pm$2.5
30% 0.0$\pm$0.0 44.3$\pm$0.7 36.0$\pm$1.1 38.7$\pm$4.3 34.3$\pm$1.9 36.7$\pm$2.4 48.0$\pm$3.1 35.7$\pm$0.3 49.3$\pm$0.7 50.7$\pm$1.4
50% 0.0$\pm$0.0 53.3$\pm$1.2 31.0$\pm$0.6 47.0$\pm$0.6 34.3$\pm$1.3 12.7$\pm$5.2 42.7$\pm$2.0 34.0$\pm$0.6 49.3$\pm$0.3 56.7$\pm$0.9
70% 0.0$\pm$0.0 39.0$\pm$1.5 30.3$\pm$0.9 46.7$\pm$1.3 36.3$\pm$0.7 4.3$\pm$3.5 51.0$\pm$1.7 33.3$\pm$0.7 31.3$\pm$0.9 52.0$\pm$1.0
90% 0.0$\pm$0.0 35.0$\pm$0.6 39.7$\pm$0.3 35.3$\pm$1.9 0.0$\pm$0.0 0.0$\pm$0.0 41.0$\pm$1.7 16.7$\pm$0.9 28.0$\pm$0.6 44.3$\pm$1.4
Re-mapped label tasks AG News Remap 0% 7.0$\pm$0.6 7.0$\pm$0.0 12.3$\pm$0.7 30.3$\pm$1.4 28.3$\pm$1.3 4.3$\pm$1.1 9.0$\pm$2.5 11.3$\pm$0.3 19.3$\pm$0.7 30.3$\pm$1.2
10% 2.0$\pm$0.0 7.0$\pm$0.0 13.0$\pm$1.1 17.0$\pm$4.0 19.0$\pm$4.5 6.0$\pm$0.8 13.0$\pm$3.2 12.0$\pm$0.6 30.0$\pm$0.6 30.3$\pm$4.4
30% 3.7$\pm$0.3 1.0$\pm$0.0 4.7$\pm$0.3 28.0$\pm$2.0 22.3$\pm$0.7 7.7$\pm$2.2 5.0$\pm$0.6 12.3$\pm$0.3 28.0$\pm$5.0 40.0$\pm$1.0
50% 4.7$\pm$0.3 3.0$\pm$0.0 3.7$\pm$0.3 27.3$\pm$2.3 21.3$\pm$2.7 8.7$\pm$3.1 13.0$\pm$3.5 6.7$\pm$0.3 21.7$\pm$2.3 40.0$\pm$3.1
70% 2.0$\pm$0.0 23.3$\pm$0.3 7.0$\pm$0.0 20.0$\pm$1.0 16.0$\pm$0.6 15.7$\pm$7.1 5.0$\pm$1.1 4.7$\pm$0.3 36.0$\pm$0.6 45.0$\pm$2.0
90% 8.0$\pm$0.6 21.0$\pm$1.1 2.0$\pm$0.0 22.7$\pm$7.0 2.0$\pm$0.0 11.7$\pm$1.9 23.0$\pm$1.5 6.3$\pm$0.3 27.3$\pm$1.9 36.7$\pm$3.2
SST5 Reverse 0% 13.3$\pm$0.7 11.7$\pm$0.7 9.7$\pm$0.9 22.0$\pm$3.6 18.3$\pm$0.7 10.3$\pm$0.3 21.7$\pm$5.5 11.3$\pm$0.3 23.7$\pm$0.7 34.0$\pm$3.0
10% 13.3$\pm$0.9 11.3$\pm$0.3 9.3$\pm$0.3 20.0$\pm$2.6 18.0$\pm$1.0 11.7$\pm$1.5 27.0$\pm$1.1 12.0$\pm$0.6 23.0$\pm$0.6 31.3$\pm$2.6
30% 15.7$\pm$0.3 13.3$\pm$0.3 11.7$\pm$0.7 19.3$\pm$2.3 23.7$\pm$0.7 9.7$\pm$0.3 21.3$\pm$2.7 11.0$\pm$0.6 13.0$\pm$1.5 24.3$\pm$1.8
50% 13.0$\pm$0.6 12.3$\pm$0.3 12.0$\pm$0.0 28.3$\pm$2.9 14.0$\pm$0.6 12.7$\pm$0.7 14.3$\pm$0.9 9.3$\pm$0.3 14.0$\pm$2.0 28.7$\pm$1.4
70% 12.0$\pm$0.0 12.0$\pm$0.6 14.3$\pm$0.3 24.7$\pm$0.7 12.0$\pm$1.1 11.7$\pm$0.5 18.7$\pm$2.9 15.7$\pm$0.3 33.0$\pm$1.0 33.3$\pm$3.8
90% 16.0$\pm$0.6 9.3$\pm$0.7 16.0$\pm$0.6 12.7$\pm$0.3 13.3$\pm$0.9 12.7$\pm$0.3 14.0$\pm$0.6 11.7$\pm$0.7 10.7$\pm$0.3 19.0$\pm$6.1
# best-performing tasks 0 0 0 0 0 0 2 0 1 21
@@
antonyms 84.7$\pm$0.3 82.0$\pm$0.6 -2.7 $\downarrow$
auto_categorization 34.0$\pm$2.1 48.0$\pm$7.2 14.0 $\uparrow$
@@
larger_animal 88.0$\pm$0.0 59.3$\pm$0.9 -28.7 $\downarrow$
negation 89.0$\pm$0.6 88.0$\pm$0.6 -1.0 $\downarrow$
object_counting 52.0$\pm$2.1 57.3$\pm$3.5 5.3 $\uparrow$
orthography_starts_with 69.3$\pm$0.7 72.7$\pm$1.2 3.3 $\uparrow$
rhymes 97.7$\pm$1.3 68.0$\pm$8.5 -29.7 $\downarrow$
second_word_letter 48.3$\pm$1.4 100.0$\pm$0.0 51.7 $\uparrow$
sentence_similarity 32.7$\pm$1.2 32.7$\pm$2.7 0.0 $\circ$
sentiment 89.0$\pm$0.0 93.7$\pm$0.3 4.7 $\uparrow$
@@
synonyms 13.7$\pm$0.3 14.0$\pm$1.5 0.3 $\uparrow$
taxonomy_animal 78.7$\pm$1.9 78.3$\pm$6.8 -0.3 $\downarrow$
translation_en-de 83.3$\pm$0.3 84.7$\pm$0.3 1.3 $\uparrow$
translation_en-es 84.7$\pm$0.9 89.3$\pm$0.3 4.7 $\uparrow$
translation_en-fr 85.0$\pm$1.0 88.0$\pm$0.6 3.0 $\uparrow$
word_sorting 69.7$\pm$0.3 73.3$\pm$0.7 3.7 $\uparrow$
word_unscrambling 62.0$\pm$2.1 62.0$\pm$1.5 0.0 $\circ$
LR (10% noise) 42.3$\pm$0.9 24.0$\pm$14.1 -18.3 $\downarrow$
LP-variant (10% noise) 48.0$\pm$2.5 55.0$\pm$1.0 7.0 $\uparrow$
AG News Remap (10% noise) 30.3$\pm$4.4 51.3$\pm$1.2 21.0 $\uparrow$
SST5 Reverse (10% noise) 31.3$\pm$2.6 34.0$\pm$1.0 2.7 $\uparrow$